# CoRe: Benchmarking LLMs’ Code Reasoning Capabilities through Static Analysis Tasks

### v2

@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable capabilities across a wide range of domains (openai2024gpt4technicalreport, ; geminiteam2025geminifamilyhighlycapable, ; wei2022chainofthought, ; yao2023react, ; grattafiori2024llama3herdmodels, ; xie2024resym, ; xie2025effective, ). In the area of software engineering, they are increasingly used in tasks such as program synthesis austin2021program ; wu2024repoformer , program repair (xia2024agentless, ; zhang2024autocoderover, ), and test generation (xia2023apr, ; schafer2024unittest, ). LLMs are usually prompted with high-level objectives xia2024agentless ; zhang2024autocoderover , such as identifying a buggy line, generating test cases to reach a target condition, or producing patches to fix faulty behavior.
@@
At the same time, researchers are beginning to use LLMs as static analyzers, applying them to downstream tasks such as vulnerability detection nunez2024autosafecoder , automatic debugging jiang2024ledex , and program repair (10.1145/3715004, ). These approaches, no matter whether implicitly relying on or explicitly prompting LLMs for code reasoning, assume that the model has mastered core code semantic analysis skills, which can support downstream clients with exceptional performance empirically. However, their underlying ability to reason about program semantics remains under-evaluated despite promising end-task results. Therefore, we urgently require an effective benchmarking solution to directly assess whether LLMs possess the deep semantic understanding and reasoning capabilities necessary to support complex software engineering tasks.
Limitations of Existing Benchmarks. Existing benchmarks mainly target the evaluation of LLMs upon code-centric tasks in an end-to-end fashion (chen2023diversevul, ; ding2024vulnerability, ; humaneval, ; liang2024can, ; CodeMMLU, ), without fundamentally assessing the model’s reasoning ability over program semantics. For example, program repair benchmarks such as SWE-Bench (jimenez2023swe, ) and SWE-Lancer (miserendino2025swe, ) provide buggy-fixed code pairs to evaluate patch generation correctness, yet lack fine‐grained ground-truth explanations of why each bug occurs. Other efforts that target fundamental program properties rather than downstream applications, like dynamic trace prediction (yadavally2024predictive, ; DBLP:conf/kbse/Jiang0CS024, ; ding2024semcoder, ), focus exclusively on observed runtime behaviors under specific inputs, leaving unexercised branches and other possible inputs overlooked. Consequently, they fail to provide adequate evaluation of whether LLMs possess the semantic reasoning required for advanced coding tasks, which highlights the need for a benchmark that directly evaluates core program analysis skills.
@@ 2 Background and Motivation
Understanding program semantics requires reasoning beyond surface-level syntax. In traditional program analysis, dependencies are central to modeling how specific program values flow between different program lines, facilitating a wide range of applications, including bug detection(Pardiff, ; FindBugs, ; 8812075, ; 10.1145/3576915.3616614, ), program optimization (10.1145/872726.806984, ; 10.1145/115372.115320, ; 10.1016/j.cose.2024.103704, ; 5185392, ), and testing(Randoop, ; AFLGO, ; Angora, ; FuzzInMem, ; 10.1145/3533767.3534403, ; AFLPlusPlus, ). In this work, we evaluate LLMs on three core reasoning tasks: data dependency, control dependency, and information flow. The first two represent fundamental program dependencies, while the third captures higher-level semantic reasoning about how program values propagate through execution. Evaluating LLMs on them provides insight into their ability to capture deep semantic structure in code.
@@ 2.1 Data Dependency
A data dependency occurs when the value of one variable depends on the value of another, typically arising when variables are assigned and then subsequently used alfred2007compilers .
@@
Applications. Data dependencies essentially induce the def-use chains in the program, serving as the ingredient for diverse static analysis applications, such as program slicing (yadavally2024predictive, ), compiler optimizations (ferrante1987program, ), and taint tracking in security (Angora, ; Arzt14FlowDroid, ), with growing interest in using LLMs to assist such analyses (wang2024llmdfa, ; li2025iris, ).
@@ 2.2 Control Dependency
Control dependency captures whether the execution of one statement is governed by another (podgurski1990formal, ). We use the line numbers of the statements to indicate the dependency. Specifically, a statement at $\ell_{2}$ is control-dependent on $\ell_{1}$ if the condition at $\ell_{1}$ determines whether $\ell_{2}$ executes. This holds when there is at least one branch from $\ell_{1}$ where $\ell_{2}$ always executes, and another where it may not.
@@
Applications. Control dependencies capture the behavior of branches. They are widely used in path reachability reasoning in various program analysis tasks, such as bug detection 10.1145/3319535.3363225 ; 10.1145/3597926.3598142 and program verification 766902 ; 10.1145/3238147.3238218 , which involves reasoning about the variables affecting the path execution.
Fig. 1b is an example of applying control dependency analysis in fuzzing, which presents a real-world vulnerability CVE-2022-26129 in the BABEL routing daemon. The call to memcpy at $\ell_{12}$ is the vulnerable operation that may cause a buffer overread. Reaching this line requires the input to satisfy a chain of conditions from $\ell_{2}$, $\ell_{4}$, $\ell_{7}$, $\ell_{9}$, and $\ell_{10}$. These conditionals form a transitive control dependency trace ending at $\ell_{12}$. Because triggering the vulnerable memory access depends on passing multiple branching checks, control dependency analysis is crucial for guiding fuzzers to explore such paths and uncover deep bugs huang2020pangolin ; jiang2023evaluating .
@@ 2.3 Information Flow
Information flow (denning1977certification, ; genaim2005information, ) captures how the value of one variable can influence another through data or control dependencies. It may be explicit, as in direct assignments, or implicit, when control conditions determine which value a variable receives. An information flow may be a sequence of implicit, explicit, or mixed flows.
@@
Applications. Information flow analysis is widely used in security and privacy, including taint tracking, non-interference, and enforcement of confidentiality and integrity policies 1159651 . In software engineering, it supports the program slicing and more downstream tasks, such as program debugging and fault localization. For example, it helps isolate the minimal subprogram relevant to a fault, improving code understanding and the developer’s ability to identify and fix bugs (francel2001value, ; yadavally2024learning, ). Recent research increasingly explores LLMs for assisting such analyses (mashnoor2025llminfoflow, ).
@@ 3.1 Benchmark Construction
We sample programs from two major sources: CodeNet (puri2021codenet, ), a large-scale dataset of competitive code, and Google Code Jam (GCJ) (gcj, ), which contains expert-written solutions to algorithmic problems. From these, we select 60 files each in C/C++, Java, and Python, resulting in 180 functions in total. We also select one target function from each file.
Our sampling ensures diverse program structures and a balanced distribution of lines of code (LoC) in each programming language. Specifically, we categorize candidate functions into four LoC buckets: 21–40, 41–60, 61–80, and 81–100, and ensure equal distribution among them. As function calls can introduce long dependency chains across functions, making it difficult to provide all the relevant functions in a single prompt, we only focus on intra-procedural analysis, excluding inter-procedural reasoning. We only select the functions that do not invoke any non-library functions, which is also a common practice in existing studies (yadavally2024predictive, ; yadavally2024learning, ), Check appendix A for benchmark construction details.
@@ 6.1 LLMs for Software Engineering Tasks
LLMs have been widely applied across many domains DBLP:conf/nips/BrownMRSKDNSSAA20 ; DBLP:journals/corr/abs-2407-15351 ; DBLP:conf/iclr/WangXZLCHXSX0LL24 ; DBLP:journals/pacmpl/Beurer-Kellner023 , including software engineering, where they are used for tasks such as programrepair (xia2024agentless, ; zhang2024autocoderover, ; xia2023apr, ; 10.1145/3658644.3691412, ), code generation jiang2024ledex ; liang2024can ; lyu2024automatic ; pei2023exploiting ; gao2025pr2 , software testing (schafer2024unittest, ; yang2024enhancing, ; zheng2025validating, ; zheng2025large, ; zheng2025llm, ), and vulnerability detection (chen2023diversevul, ; li2025iris, ; DBLP:journals/corr/abs-2406-11147, ). A key enabler for these applications is reasoning over static program properties. Traditional static analysis tools, however, are language-specific and require extensive manual engineering. To overcome these limitations, recent works increasingly use LLMs themselves to perform static analysis as an intermediate step. LLMDFA (wang2024llmdfa, ) and E&V (hao2023ev, ) apply LLM-based data- and control-flow reasoning to detect bugs; IRIS li2025iris leverages LLMs to examine data-flow paths reported by static analyzers, effectively mitigating the false positives. CoRe offers a fine-grained benchmark to evaluate LLMs’ capabilities in reasoning about program semantics, which underlie these downstream SE tasks.
@@ 6.2 Code Reasoning Benchmark and Evaluation
Code reasoning, the ability to understand program behavior, is fundamental to downstream LLM applications such as program synthesis (huang2024code, ; bhaskar2023benchmarking, ; du2023classeval, ), program repair (jimenez2023swe, ; 10.1145/3650212.3652140, ; 10.1145/3597926.3598135, ), and bug or vulnerability detection (yildiz2025benchmarking, ; 10879492, ; chen2023diversevul, ; ding2024vulnerability, ). These tasks require models to reason over program semantics, including data and control flow. Yet, most benchmarks only measure downstream success (e.g., detecting vulnerabilities successfully) without directly probing models’ reasoning capabilities on code semantics.
Recent work has begun to explicitly evaluate internal code reasoning. CRUXEval CRUXEVALX focuses solely on Python and evaluates input-output behavior, which aligns more with dynamic analysis. In contrast, our benchmark targets static analysis skills and covers Python, Java, and C/C++. CRQBench (dinella2024crqbench, ) evaluates LLMs’ semantic reasoning with C++ questions focused on variable tracking and behavioral equivalence. However, its questions focus on high-level behaviors and are automatically generated from pull requests and review comments, making them error-prone. Similarly, (pei2023exploiting, ) evaluates finetuned LLMs on several program analysis tasks, such as memory region prediction and function signature recovery from Java programs. However, the tasks remain high-level. CodeMMLU CodeMMLU focuses on high level also focuses on knowledge-based high-level questions. REval chen2024reasoning focuses on programs’ runtime behaviors such as code coverage prediction. CodeCrash (lam2025codecrash, ) evaluates model robustness via adversarial code perturbations, but does not directly assess semantic reasoning. CoRe fills this gap by enabling fine-grained assessment of models’ fundamental code understanding capabilities, including data and control dependencies and information flow, beyond surface-level pattern recognition or input-output matching.
@@ 8 Limitations
Our benchmark focuses on relatively short (within 100 LoC) due to the difficulty of scaling automated construction (Section B) and significant manual effort in annotation (Section 3.2). However, the functions within this size range still reflect real-world practices, as well-structured codebases often keep functions concise for maintainability (e.g.,  40–50 LoC) nvidiaCodeStyle ; googleCppStyleGuideWriteShortFunctions . Nonetheless, we curate 12,553 instances with an assurance of diversity and difficulty (Section 3.3.2). While annotation is manual and may introduce errors, we alleviate this through sufficient cross-validation (Section 3.2).

### v3

@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable capabilities across a wide range of domains (Achiam et al., 2023; Team et al., 2023; Wei et al., 2022; Yao et al., 2023; Grattafiori et al., 2024b; Xie et al., 2024, 2025). In the area of software engineering, they are increasingly used in tasks such as program synthesis Austin et al. (2021); Wu et al. (2024), program repair (Xia et al., 2024; Zhang et al., 2024c), and test generation (Xia et al., 2023; Schäfer et al., 2024). LLMs are usually prompted with high-level objectives Xia et al. (2024); Zhang et al. (2024c), such as identifying a buggy line, generating test cases to reach a target condition, or producing patches to fix faulty behavior.
@@
At the same time, researchers are beginning to use LLMs as static analyzers, applying them to downstream tasks such as vulnerability detection Nunez et al. (2024), automatic debugging Jiang et al. (2024a), and program repair (Li et al., 2025a). These approaches, no matter whether implicitly relying on or explicitly prompting LLMs for code reasoning, assume that the model has mastered core code semantic analysis skills, which can support downstream clients with exceptional performance empirically. However, their underlying ability to reason about program semantics remains under-evaluated despite promising end-task results. Therefore, we urgently require an effective benchmarking solution to directly assess whether LLMs possess the deep semantic understanding and reasoning capabilities necessary to support complex software engineering tasks.
Limitations of Existing Benchmarks. Existing benchmarks mainly target the evaluation of LLMs upon code-centric tasks in an end-to-end fashion (Chen et al., 2023; Ding et al., 2024a; Chen et al., 2021; Liang et al., 2025; Nguyen et al., 2025), without fundamentally assessing the model’s reasoning ability over program semantics. For example, program repair benchmarks such as SWE-Bench (Jimenez et al., 2024) and SWE-Lancer (Miserendino et al., 2025) provide buggy-fixed code pairs to evaluate patch generation correctness, yet lack fine‐grained ground-truth explanations of why each bug occurs. Other efforts that target fundamental program properties rather than downstream applications, like dynamic trace prediction (Yadavally et al., 2024a; Jiang et al., 2024b; Ding et al., 2024b), focus exclusively on observed runtime behaviors under specific inputs, leaving unexercised branches and other possible inputs overlooked. Consequently, they fail to provide adequate evaluation of whether LLMs possess the semantic reasoning required for advanced coding tasks, which highlights the need for a benchmark that directly evaluates core program analysis skills.
@@ 2 Background and Motivation
Understanding program semantics requires reasoning beyond surface-level syntax. In traditional program analysis, dependencies are central to modeling how specific program values flow between different program lines, facilitating a wide range of applications, including bug detection(Zheng et al., 2024; Ayewah et al., 2008; Fan et al., 2019; Shi et al., 2023), program optimization (Chaitin, 1982; Cytron et al., 1991; Ahmed et al., 2024; Brandt and Schneider, 2009), and testing(Pacheco and Ernst, 2007; Böhme et al., 2017; Chen and Chen, 2018; Liu et al., 2024b, 2022; Fioraldi et al., 2020). In this work, we evaluate LLMs on three core reasoning tasks: data dependency, control dependency, and information flow. The first two represent fundamental program dependencies, while the third captures higher-level semantic reasoning about how program values propagate through execution. Evaluating LLMs on them provides insight into their ability to capture deep semantic structure in code.
@@ 2.1 Data Dependency
A data dependency occurs when the value of one variable depends on the value of another, typically arising when variables are assigned and then subsequently used Alfred et al. (2007).
@@
Applications. Data dependencies essentially induce the def-use chains in the program, serving as the ingredient for diverse static analysis applications, such as program slicing (Yadavally et al., 2024a), compiler optimizations (Ferrante et al., 1987), and taint tracking in security (Chen and Chen, 2018; Arzt et al., 2014), with growing interest in using LLMs to assist such analyses (Wang et al., 2024a; Li et al., 2025b).
@@ 2.2 Control Dependency
Control dependency captures whether the execution of one statement is governed by another (Podgurski and Clarke, 1990). We use the line numbers of the statements to indicate the dependency. Specifically, a statement at $\ell_{2}$ is control-dependent on $\ell_{1}$ if the condition at $\ell_{1}$ determines whether $\ell_{2}$ executes. This holds when there is at least one branch from $\ell_{1}$ where $\ell_{2}$ always executes, and another where it may not.
@@
Applications. Control dependencies capture the behavior of branches. They are widely used in path reachability reasoning in various program analysis tasks, such as bug detection Chen et al. (2019); Zhang et al. (2023) and program verification Jensen et al. (1999); Chen and He (2018), which involves reasoning about the variables affecting the path execution.
Fig. 1b is an example of applying control dependency analysis in fuzzing, which presents a real-world vulnerability CVE-2022-26129 in the BABEL routing daemon. The call to memcpy at $\ell_{12}$ is the vulnerable operation that may cause a buffer overread. Reaching this line requires the input to satisfy a chain of conditions from $\ell_{2}$, $\ell_{4}$, $\ell_{7}$, $\ell_{9}$, and $\ell_{10}$. These conditionals form a transitive control dependency trace ending at $\ell_{12}$. Because triggering the vulnerable memory access depends on passing multiple branching checks, control dependency analysis is crucial for guiding fuzzers to explore such paths and uncover deep bugs Huang et al. (2020); Jiang et al. (2023).
@@ 2.3 Information Flow
Information flow (Denning and Denning, 1977; Genaim and Spoto, 2005) captures how the value of one variable can influence another through data or control dependencies. It may be explicit, as in direct assignments, or implicit, when control conditions determine which value a variable receives. An information flow may be a sequence of implicit, explicit, or mixed flows.
@@
Applications. Information flow analysis is widely used in security and privacy, including taint tracking, non-interference, and enforcement of confidentiality and integrity policies Sabelfeld and Myers (2003). In software engineering, it supports the program slicing and more downstream tasks, such as program debugging and fault localization. For example, it helps isolate the minimal subprogram relevant to a fault, improving code understanding and the developer’s ability to identify and fix bugs (Francel and Rugaber, 2001; Yadavally et al., 2024b). Recent research increasingly explores LLMs for assisting such analyses (Mashnoor et al., 2025).
@@ 3.1 Benchmark Construction
We sample programs from two major sources: CodeNet (Puri et al., 2021), a large-scale dataset of competitive code, and Google Code Jam (GCJ) (40), which contains expert-written solutions to algorithmic problems. From these, we select 60 files each in C/C++, Java, and Python, resulting in 180 functions in total. We also select one target function from each file.
Our sampling ensures diverse program structures and a balanced distribution of lines of code (LoC) in each programming language. Specifically, we categorize candidate functions into four LoC buckets: 21–40, 41–60, 61–80, and 81–100, and ensure equal distribution among them. As function calls can introduce long dependency chains across functions, making it difficult to provide all the relevant functions in a single prompt, we only focus on intra-procedural analysis, excluding inter-procedural reasoning. We only select the functions that do not invoke any non-library functions, which is also a common practice in existing studies (Yadavally et al., 2024a, b), Check appendix A for benchmark construction details.
@@ 6.1 LLMs for Software Engineering Tasks
LLMs have been widely applied across many domains Brown et al. (2020); Zhang et al. (2024b); Wang et al. (2024b); Beurer-Kellner et al. (2023), including software engineering, where they are used for tasks such as programrepair (Xia et al., 2024; Zhang et al., 2024c; Xia et al., 2023; Zhang et al., 2024a), code generation Jiang et al. (2024a); Liang et al. (2025); Lyu et al. (2024); Pei et al. (2024); Gao et al. (2025), software testing (Schäfer et al., 2024; Yang et al., 2024; Zheng et al., 2025b, c, a), and vulnerability detection (Chen et al., 2023; Li et al., 2025b; Du et al., 2024). A key enabler for these applications is reasoning over static program properties. Traditional static analysis tools, however, are language-specific and require extensive manual engineering. To overcome these limitations, recent works increasingly use LLMs themselves to perform static analysis as an intermediate step. LLMDFA (Wang et al., 2024a) and E&V (Hao et al., 2023) apply LLM-based data- and control-flow reasoning to detect bugs; IRIS Li et al. (2025b) leverages LLMs to examine data-flow paths reported by static analyzers, effectively mitigating the false positives. CoRe offers a fine-grained benchmark to evaluate LLMs’ capabilities in reasoning about program semantics, which underlie these downstream SE tasks.
@@ 6.2 Code Reasoning Benchmark and Evaluation
Code reasoning, the ability to understand program behavior, is fundamental to downstream LLM applications such as program synthesis (Huang et al., 2024; Bhaskar et al., 2023; Du et al., 2023), program repair (Jimenez et al., 2024; Ouyang et al., 2024; Wu et al., 2023), and bug or vulnerability detection (Yildiz et al., 2025; Tamberg and Bahsi, 2025; Chen et al., 2023; Ding et al., 2024a). These tasks require models to reason over program semantics, including data and control flow. Yet, most benchmarks only measure downstream success (e.g., detecting vulnerabilities successfully) without directly probing models’ reasoning capabilities on code semantics.
Recent work has begun to explicitly evaluate internal code reasoning. CRUXEval Xu et al. (2025) focuses solely on Python and evaluates input-output behavior, which aligns more with dynamic analysis. In contrast, our benchmark targets static analysis skills and covers Python, Java, and C/C++. CRQBench (Dinella et al., 2024) evaluates LLMs’ semantic reasoning with C++ questions focused on variable tracking and behavioral equivalence. However, its questions focus on high-level behaviors and are automatically generated from pull requests and review comments, making them error-prone. Similarly, (Pei et al., 2024) evaluates finetuned LLMs on several program analysis tasks, such as memory region prediction and function signature recovery from Java programs. However, the tasks remain high-level. CodeMMLU Nguyen et al. (2025) focuses on high level also focuses on knowledge-based high-level questions. REval Chen et al. (2025) focuses on programs’ runtime behaviors such as code coverage prediction. CodeCrash (Lam et al., 2025) evaluates model robustness via adversarial code perturbations, but does not directly assess semantic reasoning. CoRe fills this gap by enabling fine-grained assessment of models’ fundamental code understanding capabilities, including data and control dependencies and information flow, beyond surface-level pattern recognition or input-output matching.
@@ 8 Limitations
Our benchmark focuses on relatively short (within 100 LoC) due to the difficulty of scaling automated construction (Section B) and significant manual effort in annotation (Section 3.2). However, the functions within this size range still reflect real-world practices, as well-structured codebases often keep functions concise for maintainability (e.g.,  40–50 LoC) Nvidia (2025); Google (2025). Nonetheless, we curate 12,553 instances with an assurance of diversity and difficulty (Section 3.3.2). While annotation is manual and may introduce errors, we alleviate this through sufficient cross-validation (Section 3.2).